OpenStudy (anonymous):

f(x)=3x^2-2x-1 find equation of the line tangent parallel to line y=4x-8

OpenStudy (anonymous):

is the question to find a line tangent to f(x) that is also parallel to the line y = 4x - 8?

OpenStudy (anonymous):

if yeah f'(x) = 6x - 2 4 = 4x - 2 x = 3/2 y = 3(3/2)^2 - 2(3/2) - 1 y = 27 / 4 - 4 = 11 / 4 y = mx + b 11/4 = 4(3/2) + b b = 11/4 - 6 b = -13/4 y = 4x - 13/4